# Implicit Safe Set Algorithm for Provably Safe Reinforcement Learning

### v1

@@ Abstract
Deep reinforcement learning (DRL) has demonstrated impressive performance in many continuous control tasks. However, one major stumbling block to the real-world application of DRL is the lack of safety guarantees. Although DRL agents can statisfy the system safety in expectation through reward shaping, it is quite challenging to design the DRL agent to consistently meet hard constraints (e.g., safety specification) at every time step. On the other hand, existing works in the field of safe control provide guarantees on the persistent satisfaction of hard safety constraints. However, the explicit analytical system dynamics models are required in order to synthesize the safe control, and the dynamics models are typically not accessible in DRL settings. In this paper, we present a model-free safe control algorithm, implicit safe set algorithm, for synthesizing safeguards for DRL agents that will assure provable safety throughout training. The proposed algorithm synthesizes a safety index (also called the barrier certificate) and a subsequent safe control law only by querying a black-box dynamic function (e.g., a digital twin simulator). Moreover, we theoretically prove that the implicit safe set algorithm guarantees finite time convergence to the safe set and forward invariance for both continuous-time and discrete-time systems. We validate the proposed implicit safe set algorithm on the state-of-the-art safety benchmark Safety Gym, where the proposed method achieves zero safety violations and gains $95\%\pm 9\%$ cumulative reward compared to state-of-the-art safe DRL methods. Furthermore, the resulting algorithm scales well to high-dimensional systems with parallel computing.
1 Introduction
@@
Ensuring safety in terms of persistently satisfying a hard state constraint has long been considered to be an important capability for robotics (?, ?, ?). For example, vehicles should not crash into pedestrians; robot arms should not hit walls (?, ?); and aerial drones should not collide with buildings. When robots learn skills using reinforcement learning (?, ?), they need to sufficiently explore the environment (?) in order to learn an optimal policy. However, not all explorations are safe. To ensure robots learn a safe policy, the widely adopted methods are to add penalty for visiting unsafe states. However, it is not enough to simply use these posterior measures (?, ?, ?) to make the robot learn to stay away from the dangerous states. Prior measures are needed to prevent the robots from entering unsafe states in the first place.
@@
Constraining the robot motion to persistently satisfy a hard safety specification (?) in uncertain environments is an active research topic in the safe control community. The most widely used ones are energy function-based methods. These methods (?, ?, ?, ?, ?, ?) define an energy function which assigns low energy for the safe states and constructs the safe control law to dissipate energy. Then the system will always stay within the safe set (i.e., forward invariance) (?), and the system will eventually converge to safe states if it starts from an unsafe state (i.e., finite time convergence). These methods were first introduced for deterministic systems, and then have been recently extended to stochastic and uncertain systems (?, ?) (uncertainties in either the state measurements or the future evolution of the ego robot or obstacles) and combined with robot learning for safety assurance (?). However, a key drawback of these methods is that they need white-box analytical models of the system dynamics (e.g., Kinematic Bicycle Model (?)) and/or related uncertainties, which are difficult to access during robot learning.
@@
This paper extends the energy function-based methods and proposes a model-free safe control method that provides safety guarantee for the robot without the requirements of analytical dynamics model. The proposed model-free safe control is able to safeguard any robot learning algorithm. The key insight we have is that the safe control law can be synthesized without the knowledge of white-box analytical dynamics model, as long as we have a black-box dynamics function, i.e., a function that maps the current state and control to the next state. We do highlight that black-box dynamics functions are usually available in real-world applications, e.g., high-fidelity digital twin simulators (?), or deep neural network dynamics models constructed by data-driven learning (?, ?).
@@
The key questions that we want to answer are: 1) how to synthesize the energy function (in our case called the safety index) with black-box dynamics, so that there always exists a feasible control to dissipate the system energy at all possible states; 2) how to efficiently synthesize the optimal safe control with black-box dynamics, such that the control input can both dissipate the system energy and achieve good task efficiency; and 3) how to formally prove that the synthesized energy function and the generated safe control can guarantee forward invariance and finite time convergence to the safe set. We propose a safety index design rule to address the first question, and a sample-efficient algorithm to perform black-box constrained optimization to address the second question. As for the third question, we show that under certain assumptions, the proposed safety index design rule together with the proposed black-box optimization algorithm can guarantee forward invariance and finite time convergence to the safe set. By combining these approaches with the (model-based) safe set algorithm (?), we propose the (model-free) implicit safe set algorithm (ISSA). We then use ISSA to safeguard deep reinforcement learning (DRL) agents.
@@
A short version of this paper has been presented earlier (?), where ISSA is proposed and evaluated with the assumption that the sampling time is almost zero (i.e., the system evolves almost like a continuous time system). In reality, most control and simulation systems are implemented in a discrete-time manner with non-negligible sampling time. As a result, the controller may not be able to respond immediately to potential violation of safety. Therefore, assuring provable safety for discrete-time systems with non-negligible sampling time is important. In this paper, we introduce a general version of the implicit safe set algorithm which explicitly takes the sampling time into consideration. In particular, a novel convergence trigger (CTrigger) algorithm is introduced. The key contributions of this paper are summarized below:
@@
  • We propose two techniques to synthesize safe control laws without white-box models: continuous-time and discrete-time safety index design rules for mobile robots in a 2D plane, and a sample-efficient black-box optimization algorithm using adaptive momentum boundary approximation (AdamBA).
@@
  • We propose the implicit safe set algorithm (ISSA) using these two techniques, which guarantees to generate safe controls for all system states without knowing the explicit system dynamics. We show that ISSA can safeguard DRL agents to ensure zero safety violation during training in Safety Gym. Our code is available on Github. [^1] https://github.com/intelligent-control-lab/ISSA_CoRL21
@@
  • We provide the theoretical guarantees that 1) ISSA ensures the forward-invariance safety for both continuous-time and discrete-time systems; and 2) finite time convergence safety when ISSA is combined with CTrigger in for both continuous-time and discrete-time systems (i.e., with any sampling time step).
@@
In the remainder of the paper, we first formulate the mathematical problem for safeguard synthesis in Section 2. We then discuss related work about safe reinforcement and safe control in Section 3. In Section 4 and Section 5, we first introduce the implicit safe set algorithm, and provide theoretical results in continuous-time systems. We further extend the implicit safe set algorithm to the discrete-time systems and provide theoretical guarantees of forward invariance and finite time convergence in Section 6 and Section 7, respectively. Finally, we validate our proposed method in Safety Gym environments in Section 8.
2 Problem Formulation
@@
$\begin{split}&x_{t+1}=f(x_{t},u_{t}),\\ \end{split}$ (1)
where $f:\mathcal{X}\times\mathcal{U}\rightarrow\mathcal{X}$ is a function that maps the current robot state and control to the next robot state. For simplicity, this paper considers deterministic dynamics. The proposed method can be easily extended to the stochastic case through robust safe control (?, ?), which will be left for future work. Moreover, it is assumed that we can only access an implicit black-box form of $f$, e.g., as an implicit digital twin simulator or a deep neural network model. Note that the word implicit refers to that we evaluate $f(x,u)$ without any explicit knowledge or analytical form of $f(x,u)$.
@@
The safety specification requires that the system state should be constrained in a closed subset in the state space, called the safe set $\mathcal{X}_{S}$. The safe set can be represented by the zero-sublevel set of a continuous and piecewise smooth function $\phi_{0}:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}$, i.e., $\mathcal{X}_{S}=\{x\mid\phi_{0}(x)\leq 0\}$. $\mathcal{X}_{S}$ and $\phi_{0}$ are directly specified by users. The design of $\phi_{0}$ is straightforward in most scenarios. For example, for collision avoidance, $\phi_{0}$ can be designed as the negative closest distance between the robot and environmental obstacles.
Figure 1: Notations.
@@
A robot learning controller generates the nominal control which is subject to modification by the safeguard. The learning controller aims to maximize rewards in an infinite-horizon deterministic Markov decision process (MDP). An MDP is specified by a tuple $(\mathcal{X},\mathcal{U},\gamma,r,f)$, where $r:\mathcal{X}\times\mathcal{U}\rightarrow\mathbb{R}$ is the reward function, $10\leq\gamma<$ 1 is the discount factor, and $f$ is the deterministic system dynamics defined in (1).
@@
In this paper, we are focused on collision avoidance for mobile robots in a 2D plane. The robot and the obstacles are all geometrically simplified as point-mass circles with bounded collision radius. The safety specification is defined as $\phi_{0}=\max_{i}{\phi_{0}}_{i}$, where ${\phi_{0}}_{i}=d_{min}-d_{i}$, and $d_{i}$ denotes the distance between the center point of the robot and the center point of the $i$-th obstacle (static or non-static), and the radius of both obstacle and robot are considered with $d_{min}$ such that $d_{min}\geq\text{obstacle radius}+\text{robot radius}$. Denote $v$, $a$ and $w$ as the relative velocity, relative acceleration and relative angular velocity of the robot with respect to the obstacle , respectively, as shown in fig. 1.222Note that in our definition, $w=\dot{\alpha}$, where $\alpha$ is the angle between the robot’s heading vector and the vector from the robot to the obstacle. The definition of $w$ is different from the robot angular velocity in the world frame, where $w$ is influenced by the 1) the robot’s motion around the obstacle, which changes the vector from the robot to the obstacle and 2) the robot’s self-rotation in world frame, which changes the robot’s heading. Since the obstacle is treated as a circle, we do not consider its self-rotation in the world frame. In real world systems, a robot’s self-rotation is bounded, and its motion around the obstacle is also bounded. Therefore, $w$ is bounded. The remaining discussions are based on the following assumption.
Assumption 1 (2D Collision Avoidance).
1) The state space is bounded ($v\in[0,v_{max}]$), and the relative acceleration and angular velocity are bounded and both can achieve zeros, i.e., $w\in[w_{min},w_{max}]$ for $w_{min}\leq 0\leq w_{max}$ and $a\in[a_{min},a_{max}]$ for $a_{min}\leq 0\leq a_{max}$; 2) For all possible values of $a$ and $w$, there always exists a control $u$ to realize such $a$ and $w$;
These assumptions are easy to meet in practice. The bounds in the first assumption will be directly used to synthesize safety index $\phi$ (will be introduced in Section 3). The second assumption enables us to turn the question on whether these exists a feasible control in $\mathcal{U}_{S}^{D}$ to the question on whether there exists $a$ and $w$ to decrease $\phi$.
3 Related Work
@@
Safe RL methods enforce safety through either soft safety constraints or hard safety constraints (?). For example, as for the safety specification of mobile robots, soft safety contraints limit a number of crashes per trajectory, whereas hard safety constraints require mobile robots should never crash into pedestrians (?). Typical safe RL methods for soft safety constraints include risk-sensitive safe RL (?), Lagrangian methods (?), constrained policy optimization (CPO) (?) and state-wise constrained policy optimization (SCPO) (?). These methods are able to find policies that satisfy the safety constraint in expectation, but cannot ensure all visited states are safe. The methods that are more closely related to ours are safe RL methods with hard safety constraints. Safe RL methods with hard constraints can be divided into two categories: 1) safeguard learning policies using control theories; and 2) safeguard learning policies through safety layer to encourage safe actions.
@@
For the first category, safeguard methods are proposed based on 1) Control Barrier Function (CBF), 2) Hamilton-Jacobi reachability, and 3) Lyapunov method. Richard et al. (?) propose a general safe RL framework, which combines a CBF-based safeguard with RL algorithms to guide the learning process by constraining the exploratory actions to the ones that will lead to safe future states. However, this method strongly relies on the control-affine structure of the dynamics system, which restricts its applicability to general non-control-affine unknown dynamics systems. Ferlez et al. (?) also propose a ShieldNN leveraging Control Barrier Function to design a safety filter neural network with safety guarantees. However, ShieldNN is specially designed for an environment with the kinematic bicycle model (KBM) (?) as the dynamical model, which cannot be directly applied to general problems. Besides CBF, reachability analysis is also adopted for the safeguard synthesis. Fisac et al. (?) propose a general safety framework based on Hamilton-Jacobi reachability methods that works in conjunction with an arbitrary learning algorithm. However, these methods (?, ?) still rely heavily on the explicit analytic form of system dynamics to guarantee constraint satisfaction. In addition to CBF and reachability, Lyapunov methods are used to verify stability of a known system in control, which can be used to determine whether states and actions are safe or unsafe. Berkenkamp et al. (?) combine RL with Lyapunov analysis to ensure safe operation in discretized systems. Though provable safe control can be guaranteed under some Lipschitz continuity conditions, this method still requires explicit knowledge of the system dynamics (analytical form). In summary, control theory based safeguard methods provide good safety guarantees. However, those methods rely on the assumptions of 1) explicit knowledge of dynamics model or 2) control-affine dynamics model, which are hard to be satisfied in real world robotics applications.
@@
The second category to make RL meet hard safety contraints is adding a safety layer to filter the action proposed by RL policy to satisfy the safety constraints. Dalal et al. (?) propose to learn the system dynamics directly from offline collected data, and then add a safety layer that analytically solves a control correction formulation at each state to ensure every state is safe. However, the closed-form solution relies on the assumption of the linearized dynamics model and cost function, which is not true for most dynamics systems. They also assume the set of safe control can be represented by a linearized half-space for all states, which does not hold for most of the discrete-time system (i.e., safe control may not exist for some states). Yinlam et al. (?) propose to project either the policy parameters or the action to the set induced by linearized Lyapunov constraints, which still suffer from the same linear approximation error and non-control-affine systems as in  (?) and is not able to guarantee zero-violation. In contrast, our proposed method does not require the explicit form or control-affine form of system dynamics, and our method can be guaranteed to generate safe control as long as the safe set is not a collection of singleton points, and the safe set itself can be disjoint and unconnected.
@@
Representative energy function-based methods for safe control include potential field methods (?), control barrier functions (CBF) (?), safe set algorithms (SSA) (?, ?), sliding mode algorithms (?), and a wide variety of bio-inspired algorithms (?). The first step to synthesize a safe controller is to compute the desired energy function offline such that 1) the low energy states are safe and 2) there always exists a feasible control input to dissipate the energy. SSA has introduced a rule-based approach (?, ?, ?) to synthesize the energy function as a continuous, piece-wise smooth scalar function on the system state space $\phi:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}$. And the energy function $\phi(x)$ is called a safety index in this approach. The general form of the safety index was proposed as $\phi=\phi_{0}^{*}+k_{1}\dot{\phi}_{0}+\cdots+k_{n}\phi_{0}^{(n)}$ where 1) the roots of $01+k_{1}s+\ldots+k_{n}s^{n}=$ 0 are all negative real (to ensure zero-overshooting of the original safety constraints); 2) the relative degree from $\phi_{0}^{(n)}$ to $u$ is one (to avoid singularity); and 3) $\phi_{0}^{*}$ defines the same set as $\phi_{0}$ (to nonlinear shape the gradient of $\phi$ at the boundary of the safe set). It is shown in (?) that if the control input is unbounded ($\mathcal{U}=\mathbb{R}^{n_{u}}$), then there always exist a control input that satisfies the constraint $\dot{\phi}\leq 0\text{ when }\phi=0$; and if the control input always satisfies that constraint, then the set $\{x\mid\phi(x)\leq 0\}\cap\{x\mid\phi_{0}(x)\leq 0\}$ is forward invariant. In practice, the actual control signal is computed through a quadratic projection of the nominal control $u^{r}$ to the control constraint
$u=$ $\operatorname*{arg\,min}_{u\in\mathcal{U}}\|u-u^{r}\|^{2}\text{ s.t. }\dot{\phi}\leq-\eta(\phi),$ (2)
where $\dot{\phi}\leq-\eta(\phi)$ is a general form of the constraint; $\eta:\mathbb{R}\rightarrow\mathbb{R}$ is a non-decreasing function that $\eta(0)\geq 0$. For example, in CBF, $\eta(\phi)$ is designed to be $\lambda\phi$ for some positive scalar $\lambda$. In SSA, $\eta(\phi)$ is designed to be a positive constant when $\phi\geq 0$ and $\infty$- ∞ when $\phi<0$. Note there are two major differences between this paper and the existing results presented in (?). First, this paper considers constrained control space, which then requires careful selection of the parameters in $\phi$ to make sure the control constraint $\mathcal{U}_{S}(x):=\{u\in\mathcal{U}\mid\dot{\phi}\leq-\eta(\phi)\}$ is nonempty for states that $\phi\geq 0$. Secondly and most importantly, this paper considers general black-box dynamics, while the existing work considers analytical control-affine dynamics. For analytical control-affine dynamic models, $\mathcal{U}_{S}(x)$ is essentially a half-space, and the projection (2) can be efficiently computed by calling a quadratic programming solver. However, for black-box dynamics, this constraint is challenging to quantify.
Since real-world robot systems are always controlled in discrete-time with dynamics $x_{t+1}=f(x_{t},u_{t})$, we consider the discrete-time version of the set of safe control $\mathcal{U}_{S}^{D}(x):=\{u\in\mathcal{U}\mid\phi(f(x,u))\leq\max\{\phi(x)-\eta,0\}\}$ in the remaining discussions. Notice that the continuous-time system is a special case of discrete-time system where the simulation time step (sampling time) is negligible, i.e. $dt\rightarrow 0$. Additionally, $\mathcal{U}_{S}(x)$ and $\mathcal{U}_{S}^{D}(x)$ pose the similar requirements for safe control, which require $\phi$ to decrease at next time step when $\phi\geq 0$ at current time step.
4 Implicit Safe Set Algorithm for Systems with Negligible Sampling Time

### v2

@@ Abstract.
Deep reinforcement learning (DRL) has demonstrated remarkable performance in many continuous control tasks. However, a significant obstacle to the real-world application of DRL is the lack of safety guarantees. Although DRL agents can satisfy system safety in expectation through reward shaping, designing agents to consistently meet hard constraints (e.g., safety specifications) at every time step remains a formidable challenge. In contrast, existing work in the field of safe control provides guarantees on persistent satisfaction of hard safety constraints. However, these methods require explicit analytical system dynamics models to synthesize safe control, which are typically inaccessible in DRL settings. In this paper, we present a model-free safe control algorithm, the implicit safe set algorithm, for synthesizing safeguards for DRL agents that ensure provable safety throughout training. The proposed algorithm synthesizes a safety index (barrier certificate) and a subsequent safe control law solely by querying a black-box dynamic function (e.g., a digital twin simulator). Moreover, we theoretically prove that the implicit safe set algorithm guarantees finite time convergence to the safe set and forward invariance for both continuous-time and discrete-time systems. We validate the proposed algorithm on the state-of-the-art Safety Gym benchmark, where it achieves zero safety violations while gaining $95\%\pm 9\%$ cumulative reward compared to state-of-the-art safe DRL methods. Furthermore, the resulting algorithm scales well to high-dimensional systems with parallel computing.
copyright: ccdoi: 10.1613/jair.1.xxxxxjournalvolume: 4article: 6publicationmonth: 8journalyear: 2025
1. Introduction
@@
Deep reinforcement learning (DRL) has achieved impressive results in continuous control tasks (fujimoto2025generalpurposemodelfreereinforcementlearning, 16, 18, 44, 33), but its real-world deployment is hindered by the absence of hard safety guarantees. Ensuring safety in terms of persistently satisfying hard state constraints has long been recognized as a critical requirement in robotics (zhao2024absolutestatewiseconstrainedpolicy, 45, 42). For example, vehicles must avoid colliding with pedestrians; robotic arms should not strike walls; and aerial drones must steer clear of buildings. When robots learn skills via reinforcement learning, they need to sufficiently explore the environment to discover optimal policies. However, such exploration is not inherently safe. A common strategy to encourage safe behavior is to introduce penalties for visiting unsafe states. While these posterior measures are widely adopted (papini2019smoothing, 29, 31, 42), they are often insufficient to ensure robust safety. Instead, prior measures are required to proactively prevent robots from entering unsafe states in the first place.
@@
Constraining robot motion to persistently satisfy hard safety specifications in uncertain environments is an active area of research in the safe control community. Among the most widely used approaches are energy function-based methods. These methods (khatib1986real, 21, 3, 25, 19) define an energy function that assigns low energy to safe states and construct a control law that dissipates this energy. When the energy function and the control law are properly designed, the system remains within the safe set (i.e., exhibits forward invariance) and may also converge to safe states in finite time if it starts from an unsafe condition (i.e., demonstrates finite-time convergence). Initially developed for deterministic systems, these methods have been extended to handle stochastic and uncertain systems (pandya2024multimodalsafecontrolhumanrobot, 28, 9, 7, 36), where uncertainties may arise from state measurements, the future evolution of the ego robot, or the behavior of obstacles. Furthermore, they have been integrated with robot learning frameworks to enhance safety assurance (zhang2025passivitycentricsafereinforcementlearning, 38, 12, 15). However, a key limitation of energy function-based methods is their reliance on white-box analytical models of the system dynamics (e.g., the Kinematic Bicycle Model (kong2015kinematic, 22)) for both offline construction of the energy function and online computation of the safe control signal.
@@
This paper extends the class of energy function-based methods by proposing a model-free safe control approach that provides safety guarantees without requiring an analytical dynamics model. The proposed method is capable of safeguarding any robot learning algorithm by integrating safety constraints directly into the control synthesis. The key insight underlying our approach is that a safe control law can be synthesized without access to a white-box analytical dynamics model, provided that a black-box dynamics function is available—that is, a function that maps the current state and control input to the next state. Importantly, such black-box models are often accessible in real-world applications, for example, through high-fidelity digital twin simulators (abouchakra2025realissimbridgingsimtorealgap, 1, 11, 26) or deep neural network dynamics models trained in a data-driven manner (li2025continuallearningliftingkoopman, 23, 40). While this paper does not explicitly analyze formal guarantees on the accuracy of the black-box dynamics, it focuses on developing a reliable approach to synthesize the safe control strategy using only the black-box dynamics models, which is essential when designing safety-critical learning-based control systems in practice.
@@
The key questions that we want to answer are: 1) how to synthesize the energy function (in our case called the safety index) with black-box dynamics, so that there always exists a feasible control to dissipate the system energy at all potentially unsafe states; 2) how to efficiently synthesize the optimal safe control with black-box dynamics, such that the control input can both dissipate the system energy and achieve good task efficiency; and 3) how to formally prove that the synthesized energy function and the generated safe control can guarantee forward invariance and finite time convergence to the safe set. We propose a safety index design rule to address the first question, and a sample-efficient algorithm to perform black-box constrained optimization to address the second question. As for the third question, we show that under certain assumptions, the proposed safety index design rule together with the proposed black-box optimization algorithm can guarantee forward invariance and finite time convergence to a subset within the safe set. By combining these approaches with the (model-based) safe set algorithm (liu2014control, 25), we propose the (model-free) implicit safe set algorithm (ISSA). We then use ISSA to safeguard deep reinforcement learning (DRL) agents.
@@
A short version of this paper has been presented earlier (zhao2021model, 43), where ISSA is proposed and evaluated with the assumption that the sampling time is almost zero (i.e., the system evolves almost like a continuous-time system). In reality, most control and simulation systems are implemented in a discrete-time manner with non-negligible sampling time. As a result, the controller may not be able to respond immediately to potential violations of safety. Therefore, assuring provable safety for discrete-time systems with non-negligible sampling time is important. In this paper, we introduce a general version of the implicit safe set algorithm which explicitly takes the sampling time into consideration. In particular, a novel convergence trigger (CTrigger) algorithm is introduced. The key contributions of this paper are summarized below:
@@
  • We propose two techniques to enable safe control with black-box models: (offline) continuous-time and discrete-time safety index design rules for mobile robots in a 2D plane, and a (online) sample-efficient black-box optimization algorithm using adaptive momentum boundary approximation (AdamBA).
@@
  • We propose the implicit safe set algorithm (ISSA) using these two techniques, which guarantees to generate safe controls for all system states without knowing the explicit system dynamics. We show that ISSA can safeguard DRL agents to ensure zero safety violation during training in Safety Gym. Our code is available on Github. [^1] https://github.com/intelligent-control-lab/Implicit_Safe_Set_Algorithm
@@
  • We provide the theoretical guarantees that 1) ISSA ensures the forward-invariance safety for both continuous-time and discrete-time systems; and 2) finite time convergence safety when ISSA is combined with CTrigger for both continuous-time and discrete-time systems (i.e., with any sampling time step).
@@
In the remainder of the paper, we first formulate the mathematical problem in Section 2. We then discuss related work about safe reinforcement and safe control in Section 3. In Section 4 and Section 5, we first introduce the implicit safe set algorithm, and provide theoretical results in continuous-time systems. We further extend the implicit safe set algorithm to the discrete-time systems and provide theoretical guarantees of forward invariance and finite time convergence in Section 6 and Section 7, respectively. Finally, we validate our proposed method in Safety Gym environments in Section 8.
2. Problem Formulation
@@
[…]
where $f:\mathcal{X}\times\mathcal{U}\rightarrow\mathcal{X}$ is a function that maps the current robot state and control to the next robot state. For simplicity, this paper considers deterministic dynamics. The proposed method can be easily extended to the stochastic case through robust safe control (emam2022safereinforcementlearningusing, 13), which will be left for future work. Moreover, it is assumed that we can only access an implicit black-box form of $f$, e.g., as an implicit digital twin simulator or a deep neural network model. Note that the word implicit refers to that we can only do point-wise evaluation of $f(x,u)$ without any explicit knowledge or analytical form of $f(x,u)$.
Figure 1. Visualization of robot notations.
@@
The safety specification requires that the system state should be constrained in a regular and closed subset in the state space, called the safe set $\mathcal{X}_{S}$. The safe set can be represented by the zero-sublevel set of a continuous and piecewise smooth function $\phi_{0}:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}$, i.e., $\mathcal{X}_{S}=\{x\mid\phi_{0}(x)\leq 0\}$. $\mathcal{X}_{S}$ and $\phi_{0}$ are directly specified by users. A state $x(t_{0})$ is considered safe if it belongs to the safe set $\mathcal{X}_{S}$ at time $t_{0}$, notwithstanding the possibility that the trajectory $x(t)$ may exit $\mathcal{X}_{S}$ at some later time $t_{1}>t_{0}$. The notion for the state, conditioned on the control policy, to result in a trajectory that always lies in the safe set is called invariantly safe, which is later captured by the notion of forward invariance. Hence the safe set $\mathcal{X}_{S}$ captures static safety and forward invariance, which usually happens within a subset of $\mathcal{X}_{S}$, captures dynamic safety. Nevertheless, the design of $\phi_{0}$ is straightforward in most scenarios. For example, for collision avoidance, $\phi_{0}$ can be designed as the negative closest distance between the robot and environmental obstacles.
@@
A robot learning controller generates the nominal control which is subject to modification by the safeguard. The learning controller aims to maximize rewards in an infinite-horizon deterministic Markov decision process (MDP). An MDP is specified by a tuple $(\mathcal{X},\mathcal{U},\gamma,r,f)$, where $r:\mathcal{X}\times\mathcal{U}\rightarrow\mathbb{R}$ is the reward function, $0\leq\gamma<1$ is the discount factor, and $f$ is the deterministic system dynamics defined in (1).
@@
In this paper, we are focused on collision avoidance for mobile robots in a 2D plane. The robot and the obstacles are all geometrically simplified as point-mass circles with bounded collision radius. The safety specification is defined as $\phi_{0}=\max_{i}{\phi_{0}}_{i}$, where ${\phi_{0}}_{i}=d_{min}-d_{i}$, and $d_{i}$ denotes the distance between the center point of the robot and the center point of the $i$-th obstacle (static or non-static), and the radius of both the obstacle and the robot are considered in $d_{min}$ such that $d_{min}\geq\text{obstacle radius}+\text{robot radius}$. Denote $v$, $a$ and $w$ as the relative velocity, relative acceleration and relative angular velocity of the robot with respect to the obstacle, respectively, as shown in fig. 1.222Note that in our definition, $w=\dot{\alpha}$, where $\alpha$ is the angle between the robot’s heading vector and the vector from the robot to the obstacle. The definition of $w$ is different from the robot angular velocity in the world frame, where $w$ is influenced by the 1) the robot’s motion around the obstacle, which changes the vector from the robot to the obstacle and 2) the robot’s self-rotation in world frame, which changes the robot’s heading. Since the obstacle is treated as a circle, we do not consider its self-rotation in the world frame. In real world systems, a robot’s self-rotation is bounded, and its motion around the obstacle is also bounded. Therefore, $w$ is bounded. The remaining discussions are based on the following assumption.
Assumption 1 (Actuation Reachability).
1) The state space is bounded ($v\in[0,v_{max}]$), and the relative acceleration and angular velocity are bounded and both can achieve zeros, i.e., $w\in[w_{min},w_{max}]$ for $w_{min}\leq 0\leq w_{max}$ and $a\in[a_{min},a_{max}]$ for $a_{min}\leq 0\leq a_{max}$; 2) For all possible values of $a$ and $w$, there always exists an input $u$ such that the closed-loop low-level servo tracks that pair within the sampling period $dt$.
These assumptions are easy to meet in practice. The bounds in the first assumption will be directly used to synthesize the safety index $\phi$ (will be introduced in Section 3). The second assumption enables us to turn the question on whether there exists a feasible safe control in $\mathcal{U}$ to the question on whether there exists $a$ and $w$ to influence $\phi$. The requirement is satisfied by most wheeled or tracked mobile bases whose wheel/track torques can be commanded independently. It does not imply that the robot should be able to perform holonomic motion in Cartesian space; translational motion remains subject to the usual non-holonomic constraint. Our experiments use MuJoCo’s unicycle and quadruped models whose low-level velocity controllers meet this reachability property.
3. Related Work
@@
Safe RL methods enforce safety through either soft safety constraints or hard safety constraints. For example, as for the safety specification of mobile robots, soft safety contraints limit a number of crashes per trajectory, whereas hard safety constraints require mobile robots should never crash into pedestrians. Typical safe RL methods for soft safety constraints include risk-sensitive safe RL (11039715, 27, 17), Lagrangian methods (ray2019benchmarking, 32), trust-region-based constrained policy optimization  (achiam2017constrained, 2, 41, 45). These methods are able to find policies that satisfy the safety constraint in expectation, but cannot ensure all visited states are safe. The methods that are more closely related to ours are safe RL methods with hard safety constraints. Safe RL methods with hard constraints can be divided into two categories: 1) safeguarding learning policies using control theories; and 2) safeguarding learning policies through a safety layer to encourage safe actions.
@@
For the first category, safeguard methods are proposed based on 1) Control Barrier Function (CBF), 2) Hamilton-Jacobi reachability, and 3) Lyapunov method. Richard et al. (cheng2019end, 7) propose a general safe RL framework, which combines a CBF-based safeguard with RL algorithms to guide the learning process by constraining the exploratory actions to the ones that will lead to safe future states. However, this method strongly relies on the control-affine structure of the dynamics system, which restricts its applicability to general non-control-affine unknown dynamics systems. Ferlez et al. (ferlez2020shieldnn, 14) also propose a ShieldNN leveraging Control Barrier Function to design a safety filter neural network with safety guarantees. However, ShieldNN is specially designed for an environment with the kinematic bicycle model (KBM) (kong2015kinematic, 22) as the dynamical model, which cannot be directly applied to general problems. Besides CBF, reachability analysis is also adopted for the safeguard synthesis. Fisac et al. (fisac2018general, 15) propose a general safety framework based on Hamilton-Jacobi reachability methods that works in conjunction with an arbitrary learning algorithm. However, these methods (fisac2018general, 15, 30) still rely heavily on the explicit analytic form of system dynamics to guarantee constraint satisfaction. In addition to CBF and reachability, Lyapunov methods are used to verify stability of a known system in control, which can be used to determine whether states and actions are safe or unsafe. Berkenkamp et al. (berkenkamp2017safe, 6) combine RL with Lyapunov analysis to ensure safe operation in discretized systems. Though provable safe control can be guaranteed under some Lipschitz continuity conditions, this method still requires explicit knowledge of the system dynamics (analytical form). In summary, control theory based safeguard methods provide good safety guarantees. However, those methods rely on the assumptions of 1) explicit knowledge of dynamics model or 2) control-affine dynamics model, which are hard to be satisfied in real world robotics applications.
@@
The second category to make RL meet hard safety constraints is shield synthesis–style approaches, where a safety layer monitors the action proposed by the RL policy and corrects it, if necessary, to satisfy safety constraints. Such methods can guarantee safety online by projecting unsafe actions into a certified safe set before execution. Dalal et al. (dalal2018safe, 10) propose to learn the system dynamics directly from offline collected data, and then add a safety layer that analytically solves a control correction formulation at each state to ensure every state is safe. However, the closed-form solution relies on the assumption of the linearized dynamics model and cost function, which is not true for most dynamics systems. They also assume the set of safe control can be represented by a linearized half-space for all states, which does not hold for most of the discrete-time system (i.e., safe control may not exist for some states). Yinlam et al. (chow2019lyapunov, 8) propose to project either the policy parameters or the action to the set induced by linearized Lyapunov constraints, which still suffer from the same linear approximation error and non-control-affine systems as in (dalal2018safe, 10) and is not able to guarantee zero-violation. Bejarano et al. (Bejarano_2025, 5) integrate safety filters into the RL training process, allowing the controller to adapt to the presence of the filter and thereby improving the overall safety–performance trade-off and sample efficiency. However, their approach still fundamentally relies on an analytical dynamics model, which limits applicability in scenarios where such a model is unavailable or inaccurate. These works all fall under the general paradigm of shield synthesis in RL, where a pre-specified or learned model is used online to intercept and correct unsafe actions before execution. In contrast, our proposed method does not require the explicit form or control-affine form of system dynamics, and our method can be guaranteed to generate safe control as long as the safe set is regular (i.e., the set equal to the closure of its interior; for example, it is not a collection of singleton points) and contains forward invariant subsets.
@@
[…]
(2) $u=$ $\operatorname*{arg\,min}_{u\in\mathcal{U}}\|u-u^{r}\|^{2}\text{ s.t. }\dot{\phi}\leq-\eta(\phi),$
where $\dot{\phi}\leq-\eta(\phi)$ is a general form of the constraint; $\eta:\mathbb{R}\rightarrow\mathbb{R}$ is a non-decreasing function that $\eta(0)\geq 0$. For example, in CBF, $\eta(\phi)$ is designed to be $\lambda\phi$ for some positive scalar $\lambda$. In SSA, $\eta(\phi)$ is designed to be a positive constant when $\phi\geq 0$ and -$\infty$ when $\phi<0$. Note there are two major differences between this paper and the existing results presented in (liu2014control, 25). First, this paper considers constrained control space, which then requires careful selection of the parameters in $\phi$ to make sure the control constraint $\mathcal{U}_{S}(x):=\{u\in\mathcal{U}\mid\dot{\phi}\leq-\eta(\phi)\}$ is nonempty for states that $\phi\geq 0$. Secondly and most importantly, this paper considers general black-box dynamics, while the existing work considers analytical control-affine dynamics. For analytical control-affine dynamic models, $\mathcal{U}_{S}(x)$ is essentially a half-space, and the projection (2) can be efficiently computed by calling a quadratic programming solver. However, for black-box dynamics, this constraint is challenging to quantify.
[…]
4. Implicit Safe Set Algorithm for Systems with Negligible Sampling Time